# KG-QAGen: A Knowledge-Graph-Based Framework for Systematic Question Generation and Long-Context LLM Evaluation

### v1

@@ Appendix C All templates
Table 5: Level 1 Question Templates.
Template Based On P H #SO
What is the position of [Person Name]? [if one] 0 1 0
In what organization does [Person Name] work? 0 1 0
Who is the representative of [Org Name]? [if one] 0 1 0
What is the role of [Org Name] in the agreement? [if one] 0 1 0
What company is the [Org Role (+ Sub-Role)] in the agreement? [if one] 0 1 0
What is the location of [Org Name]? [if one] 0 1 0
Which company is associated with [Location]? 0 1 0
What type of location is [Location] (e.g., Headquarters, Trade Operations, etc.)? [if one] 0 1 0
Table 6: Level 2 Question Templates.
Template Based On P H #SO
What are the positions of [Person Name]? What is the position of [Person Name]? [if one] 1 1 0
Who are the representatives of [Org Name]? Who is the representative of [Org Name]? [if one] 1 1 0
What are the roles of [Org Name] in the agreement? What is the role of [Org Name] in the agreement? [if one] 1 1 0
What companies are the [Org Role (+ Sub-Role)] in the agreement? What company is the [Org Role (+ Sub-Role)] in the agreement? [if one] 1 1 0
What are the locations of [Org Name]? What is the location of [Org Name]? [if one] 1 1 0
What types of location is [Location] (e.g., Headquarters, Trade Operations, etc.)? What type of location is [Location] (e.g., Headquarters, Trade Operations, etc.)? [if one] 1 1 0
What is the position held by both [Person Name 1] and [Person Name 2]? [if one] What is the position of [Person Name]? [if one] 0 1 1
What role do both [Org Name 1] and [Org Name 2] have in the agreement? [if one] What is the role of [Org Name] in the agreement? [if one] 0 1 1
What company is both the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] in the agreement? [if one] What company is the [Org Role (+ Sub-Role)] in the agreement? [if one] 0 1 1
Who is the [Person Position] of [Org Name]? [if one] 0 2 0
@@
What is the role in the agreement of the company where [Person Name] is employed? [if one] What is the role of [Org Name] in the agreement? [if one] 0 2 0
@@
What is the role in the agreement of the company associated with [Location]? [if one] What is the role of [Org Name] in the agreement? [if one] 0 2 0
What is the [Location Type] office of [Org Name]? 0 2 0
Table 7: Level 3 Question Templates.
Template Based On P H #SO
What are the positions held by both [Person Name 1] and [Person Name 2]? What are the positions of [Person Name]? What is the position held by both [Person Name 1] and [Person Name 2]? [if one] 1 1 1
What roles do both [Org Name 1] and [Org Name 2] have in the agreement? What are the roles of [Org Name] in the agreement? What role do both [Org Name 1] and [Org Name 2] have in the agreement? [if one] 1 1 1
What companies are both the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] in the agreement? What companies are the [Org Role (+ Sub-Role)] in the agreement? 1 1 1
What is the position held by [Person Name 1] but not by [Person Name 2]? [if one] What is the position held by both [Person Name 1] and [Person Name 2]? [if one] 0 1 2
@@
What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2]? [if one] What role do both [Org Name 1] and [Org Name 2] have in the agreement? [if one] 0 1 2
What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement? [if one] What company is both the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] in the agreement? [if one] 0 1 2
Who are the [Person Position]s of [Org Name]? Who is the [Person Position] of [Org Name]? [if one] 1 2 0
@@
What are the roles in the agreement of the company where [Person Name] is employed? What is the role in the agreement of the company where [Person Name] is employed? [if one] 1 2 0
@@
What are the roles in the agreement of the company associated with [Location]? What is the role in the agreement of the company associated with [Location]? [if one] 1 2 0
Who is the [Person Position 1] and [Person Position 2] of [Org Name]? [if one] Who is the [Person Position] of [Org Name]? [if one] 0 2 1
@@
Who is the [Person Position] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] Who is the [Person Position] of [Org Name]? [if one] 0 3 0
Who is the [Person Position] of the company associated with [Location]? [if one] Who is the [Person Position] of [Org Name]? [if one] 0 3 0
Who is the [Person Position] of the company associated where [Person Name] is employed? [if one] Who is the [Person Position] of [Org Name]? [if one] 0 3 0
@@
What is the address of [Location Type] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] What is the [Location Type] office of [Org Name]? 0 3 0
What is the address of [Location Type] of the company associated with [Location (different)]? What is the [Location Type] office of [Org Name]? 0 3 0
@@
What is the address of [Location Type] of the company where [Person Name] is employed? What is the [Location Type] office of [Org Name]? 0 3 0
@@
Template Based On P H #SO
What are the positions held by [Person Name 1] but not by [Person Name 2]? What are the positions held by both [Person Name 1] and [Person Name 2]? What is the position held by [Person Name 1] but not by [Person Name 2]? [if one] 1 1 2
What roles does [Org Name 1] have in the agreement which are not the roles of [Org Name 2]? What roles do both [Org Name 1] and [Org Name 2] have in the agreement? What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2]? [if one] 1 1 2
What companies are the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement? What companies are both the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] in the agreement? What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement? [if one] 1 1 2
What is the position held by [Person Name 1] but not by [Person Name 2] or [Person Name 3]? [if one] What is the position held by [Person Name 1] but not by [Person Name 2]? [if one] 0 1 3
What is the position held by [Person Name 1] and [Person Name 2] but not by [Person Name 3]? [if one] What is the position held by [Person Name 1] but not by [Person Name 2]? [if one] 0 1 3
What role do [Org Name 1] and [Org Name 2] have in the agreement which is not the role of [Org Name 3]? [if one] What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2]? [if one] 0 1 3
What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2] or [Org Name 3]? [if one] What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2]? [if one] 0 1 3
What company is the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] but not the [Org Role (+ Sub-Role) 3] in the agreement? [if one] What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement? [if one] 0 1 3
What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] [Org Role (+ Sub-Role) 3] in the agreement? [if one] What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement? [if one] 0 1 3
@@
Template Based On P H #SO
Who are the both [Person Position 1]s and [Person Position 2]s of [Org Name]? Who are the [Person Position]s of [Org Name]? Who is the [Person Position 1] and [Person Position 2] of [Org Name]? [if one] 1 2 1
Who are the [Person Position]s of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [the company should be uniquely identifiable] Who is the [Person Position] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] 1 3 0
Who are the [Person Position]s of the company associated with [Location]? Who is the [Person Position] of the company associated with [Location]? [if one] 1 3 0
Who are the [Person Position]s of the company associated where [Person Name] is employed? Who is the [Person Position] of the company associated where [Person Name] is employed? [if one] 1 3 0
Who is the [Person Position] of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement? [if one, and the company should be uniquely identifiable] Who is the [Person Position] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] 0 3 1
Who is both the [Person Position 1] and [Person Position 2] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] Who is the [Person Position] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] 0 3 1
Who is both the [Person Position 1] and [Person Position 2] of the company associated with [Location]? [if one] Who is the [Person Position] of the company associated with [Location]? [if one] 0 3 1
Who is both the [Person Position 1] and [Person Position 2] of the company associated where [Person Name] is employed? [if one] Who is the [Person Position] of the company associated where [Person Name] is employed? [if one] 0 3 1
What is the address of the [Location Type] office of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement? [if one, and the company should be uniquely identifiable] What is the [Location Type] office of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] 0 3 1
@@
Template Based On P H #SO
What are the positions held by [Person Name 1] but not by [Person Name 2] or [Person Name 3]? What are the positions held by [Person Name 1] but not by [Person Name 2]? What is the position held by [Person Name 1] but not by [Person Name 2] or [Person Name 3]? [if one] 1 1 3
What are the positions held by [Person Name 1] and [Person Name 2] but not by [Person Name 3]? What are the positions held by [Person Name 1] but not by [Person Name 2]? What is the position held by [Person Name 1] and [Person Name 2] but not by [Person Name 3]? [if one] 1 1 3
What roles do [Org Name 1] and [Org Name 2] have in the agreement which are not the roles of [Org Name 3]? What roles does [Org Name 1] have in the agreement which are not the roles of [Org Name 2]? What role do [Org Name 1] and [Org Name 2] have in the agreement which is not the role of [Org Name 3]? [if one] 1 1 3
What roles does [Org Name 1] have in the agreement which are not the roles of [Org Name 2] or [Org Name 3]? What roles does [Org Name 1] have in the agreement which are not the roles of [Org Name 2]? What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2] or [Org Name 3]? [if one] 1 1 3
What companies are the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] but not the [Org Role (+ Sub-Role) 3] in the agreement? What companies are the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement? What company is the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] but not the [Org Role (+ Sub-Role) 3] in the agreement? [if one] 1 1 3
What companies are the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] [Org Role (+ Sub-Role) 3] in the agreement? [if one] What companies are the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement? What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] [Org Role (+ Sub-Role) 3] in the agreement? [if one] 1 1 3
@@
Who are the [Person Position]s of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement? [the company should be uniquely identifiable] Who are the [Person Position]s of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [the company should be uniquely identifiable] Who is the [Person Position] of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement? [if one, and the company should be uniquely identifiable] 1 3 1
@@
Who are both the [Person Position 1] and [Person Position 2] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [the company should be uniquely identifiable] Who are the [Person Position]s of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [the company should be uniquely identifiable] Who is both the [Person Position 1] and [Person Position 2] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] 1 3 1
Who are both the [Person Position 1] and [Person Position 2] of the company associated with [Location]? Who are the [Person Position]s of the company associated with [Location]? Who is both the [Person Position 1] and [Person Position 2] of the company associated with [Location]? [if one] 1 3 1
@@
Who are both the [Person Position 1] and [Person Position 2] of the company associated where [Person Name] is employed? Who are the [Person Position]s of the company associated where [Person Name] is employed? Who is both the [Person Position 1] and [Person Position 2] of the company associated where [Person Name] is employed? [if one] 1 3 1
@@
Who is the [Person Position 1] but not [Person Position 2] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] Who is both the [Person Position 1] and [Person Position 2] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] 0 3 2
Who is the [Person Position 1] but not [Person Position 2] of the company associated with [Location]? [if one] Who is both the [Person Position 1] and [Person Position 2] of the company associated with [Location]? [if one] 0 3 2
@@
Who is the [Person Position 1] but not [Person Position 2] of the company associated where [Person Name] is employed? [if one] Who is both the [Person Position 1] and [Person Position 2] of the company associated where [Person Name] is employed? [if one] 0 3 2
@@
What is the [Location Type] office of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] but not the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement? [if one, and the company should be uniquely identifiable] What is the [Location Type] office of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement? [if one, and the company should be uniquely identifiable] 0 3 2

### v2

@@ Appendix B Knowledge Graph Construction Details
B.1 Qualitative Analysis of the Constructed KGs
After generating question-answer pairs from the extracted graphs, we observed two common issues affecting the quality and consistency of the QAs.
  • Duplicate entries: Multiple questions or answers representing the same entity or relation, differing only by capitalization, whitespace, punctuation, or pluralization.
  • Noisy text: URL-encoded characters (e.g., %26, %20) and miscellaneous special characters were present in questions and answers.
To address these issues, we applied a structured cleaning procedure after QA generation, ensuring the graphs themselves were not altered but the extracted QAs were reliable:
  • Text normalization:
    • Convert all text to lowercase.
    • Remove URL-encoded sequences and special characters, preserving commas for multiple answers.
    • Collapse multiple whitespace characters.
  • Duplicate answer removal:
    • Split multi-answer entries by commas and clean each entry.
    • Normalize text (lowercase, remove spaces) and reduce plural forms to singular where appropriate.
    • Remove duplicates while preserving the first occurrence.
    • Log all removed duplicates, including document identifiers, for traceability.
@@ Appendix C All templates
This appendix contains the defined question templates.
Table 5: Level 1 Question Templates with Counts.
Template Based On P H #SO Count
What is the position of [Person Name]? [if one] 0 1 0 1252
In what organization does [Person Name] work? 0 1 0 1288
Who is the representative of [Org Name]? [if one] 0 1 0 1358
What is the role of [Org Name] in the agreement? [if one] 0 1 0 845
What company is the [Org Role (+ Sub-Role)] in the agreement? [if one] 0 1 0 444
What is the location of [Org Name]? [if one] 0 1 0 790
Which company is associated with [Location]? 0 1 0 530
What type of location is [Location] (e.g., Headquarters, Trade Operations, etc.)? [if one] 0 1 0 43
Table 6: Level 2 Question Templates with Counts.
Template Based On P H #SO Count
Who is the [Person Position] of [Org Name]? [if one] 0 2 0 2274
What is the position held by both [Person Name 1] and [Person Name 2]? [if one] What is the position of [Person Name]? [if one] 0 1 1 1612
@@
What is the role in the agreement of the company where [Person Name] is employed? [if one] What is the role of [Org Name] in the agreement? [if one] 0 2 0 613
What are the roles of [Org Name] in the agreement? What is the role of [Org Name] in the agreement? [if one] 1 1 0 409
What companies are the [Org Role (+ Sub-Role)] in the agreement? What company is the [Org Role (+ Sub-Role)] in the agreement? [if one] 1 1 0 304
What role do both [Org Name 1] and [Org Name 2] have in the agreement? [if one] What is the role of [Org Name] in the agreement? [if one] 0 1 1 257
@@
What is the role in the agreement of the company associated with [Location]? [if one] What is the role of [Org Name] in the agreement? [if one] 0 2 0 180
Who are the representatives of [Org Name]? Who is the representative of [Org Name]? [if one] 1 1 0 160
What are the positions of [Person Name]? What is the position of [Person Name]? [if one] 1 1 0 136
What are the locations of [Org Name]? What is the location of [Org Name]? [if one] 1 1 0 91
In what organizations does [Person Name] work? In what organization does [Person Name] work? 1 1 0 90
What is the [Location Type] office of [Org Name]? 0 2 0 70
What types of location is [Location] (e.g., Headquarters, Trade Operations, etc.)? What type of location is [Location] (e.g., Headquarters, Trade Operations, etc.)? [if one] 1 1 0 11
Table 7: Level 3 Question Templates with Counts.
Template Based On P H #SO Count
What are the positions held by both [Person Name 1] and [Person Name 2]? What are the positions of [Person Name]? 1 1 1 3
What is the position held by [Person Name 1] but not by [Person Name 2]? [if one] What are the positions held by both [Person Name 1] and [Person Name 2]? 0 1 2 8
What roles do both [Org Name 1] and [Org Name 2] have in the agreement? What are the roles of [Org Name] in the agreement? 1 1 1 92
@@
What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2]? [if one] What roles do both [Org Name 1] and [Org Name 2] have in the agreement? 0 1 2 278
What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement? [if one] What companies are both the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] in the agreement? 0 1 2 56
Who are the [Person Position]s of [Org Name]? Who is the [Person Position] of [Org Name]? [if one] 1 2 0 57
Who is the [Person Position 1] and [Person Position 2] of [Org Name]? [if one] Who is the [Person Position] of [Org Name]? [if one] 0 2 1 1346
@@
What are the roles in the agreement of the company where [Person Name] is employed? What is the role in the agreement of the company where [Person Name] is employed? [if one] 1 2 0 318
@@
What are the roles in the agreement of the company associated with [Location]? What is the role in the agreement of the company associated with [Location]? [if one] 1 2 0 169
@@
Who is the [Person Position] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] Who is the [Person Position] of [Org Name]? [if one] 0 3 0 262
Who is the [Person Position] of the company associated with [Location]? [if one] Who is the [Person Position] of [Org Name]? [if one] 0 3 0 616
Who is the [Person Position] of the company where [Person Name] is employed? [if one] Who is the [Person Position] of [Org Name]? [if one] 0 3 0 289
@@
What is the address of [Location Type] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] What is the [Location Type] office of [Org Name]? 0 3 0 13
@@
What is the address of [Location Type] of the company where [Person Name] is employed? What is the [Location Type] office of [Org Name]? 0 3 0 75
Who is the [Person Position] of the company where [Person Name] is employed? [if one] Who is the [Person Position] of [Org Name]? [if one] 0 3 0 289
@@
Template Based On P H #SO Count
What are the positions held by [Person Name 1] but not by [Person Name 2]? What are the positions held by both [Person Name 1] and [Person Name 2]? What is the position held by [Person Name 1] but not by [Person Name 2]? [if one] 1 1 2 2
What roles does [Org Name 1] have in the agreement which are not the roles of [Org Name 2]? What roles do both [Org Name 1] and [Org Name 2] have in the agreement? What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2]? [if one] 1 1 2 134
What companies are the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement? What companies are both the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] in the agreement? What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement? [if one] 1 1 2 124
What is the position held by [Person Name 1] but not by [Person Name 2] or [Person Name 3]? [if one] What is the position held by [Person Name 1] but not by [Person Name 2]? [if one] 0 1 3 50
What is the position held by [Person Name 1] and [Person Name 2] but not by [Person Name 3]? [if one] What is the position held by [Person Name 1] but not by [Person Name 2]? [if one] 0 1 3 4
What role do [Org Name 1] and [Org Name 2] have in the agreement which is not the role of [Org Name 3]? [if one] What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2]? [if one] 0 1 3 0
What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2] or [Org Name 3]? [if one] What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2]? [if one] 0 1 3 0
What company is the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] but not the [Org Role (+ Sub-Role) 3] in the agreement? [if one] What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement? [if one] 0 1 3 1553
What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] [Org Role (+ Sub-Role) 3] in the agreement? [if one] What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement? [if one] 0 1 3 816
@@
Template Based On P H #SO Count
Who are the both [Person Position 1]s and [Person Position 2]s of [Org Name]? Who are the [Person Position]s of [Org Name]? Who is the [Person Position 1] and [Person Position 2] of [Org Name]? [if one] 1 2 1 0
Who are the [Person Position]s of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [the company should be uniquely identifiable] Who is the [Person Position] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] 1 3 0 40
Who are the [Person Position]s of the company associated with [Location]? Who is the [Person Position] of the company associated with [Location]? [if one] 1 3 0 22
Who are the [Person Position]s of the company associated where [Person Name] is employed? Who is the [Person Position] of the company associated where [Person Name] is employed? [if one] 1 3 0 5
Who is the [Person Position] of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement? [if one, and the company should be uniquely identifiable] Who is the [Person Position] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] 0 3 1 187
Who is both the [Person Position 1] and [Person Position 2] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] Who is the [Person Position] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] 0 3 1 61
Who is both the [Person Position 1] and [Person Position 2] of the company associated with [Location]? [if one] Who is the [Person Position] of the company associated with [Location]? [if one] 0 3 1 59
Who is both the [Person Position 1] and [Person Position 2] of the company associated where [Person Name] is employed? [if one] Who is the [Person Position] of the company associated where [Person Name] is employed? [if one] 0 3 1 23
What is the address of the [Location Type] office of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement? [if one, and the company should be uniquely identifiable] What is the [Location Type] office of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] 0 3 1 14
@@
Template Based On P H #SO Count
What are the positions held by [Person Name 1] but not by [Person Name 2] or [Person Name 3]? What are the positions held by [Person Name 1] but not by [Person Name 2]? What is the position held by [Person Name 1] but not by [Person Name 2] or [Person Name 3]? [if one] 1 1 3 8
What are the positions held by [Person Name 1] and [Person Name 2] but not by [Person Name 3]? What are the positions held by [Person Name 1] but not by [Person Name 2]? What is the position held by [Person Name 1] and [Person Name 2] but not by [Person Name 3]? [if one] 1 1 3 49
What roles do [Org Name 1] and [Org Name 2] have in the agreement which are not the roles of [Org Name 3]? What roles does [Org Name 1] have in the agreement which are not the roles of [Org Name 2]? What role do [Org Name 1] and [Org Name 2] have in the agreement which is not the role of [Org Name 3]? [if one] 1 1 3 0
What roles does [Org Name 1] have in the agreement which are not the roles of [Org Name 2] or [Org Name 3]? What roles does [Org Name 1] have in the agreement which are not the roles of [Org Name 2]? What role does [Org Name 1] have in the agreement which is not the role of [Org Name 2] or [Org Name 3]? [if one] 1 1 3 0
What companies are the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] but not the [Org Role (+ Sub-Role) 3] in the agreement? What companies are the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement? What company is the [Org Role (+ Sub-Role) 1] and [Org Role (+ Sub-Role) 2] but not the [Org Role (+ Sub-Role) 3] in the agreement? [if one] 1 1 3 122
What companies are the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] [Org Role (+ Sub-Role) 3] in the agreement? [if one] What companies are the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] in the agreement? What company is the [Org Role (+ Sub-Role) 1] but not the [Org Role (+ Sub-Role) 2] [Org Role (+ Sub-Role) 3] in the agreement? [if one] 1 1 3 0
@@
Who are the [Person Position]s of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement? [the company should be uniquely identifiable] Who are the [Person Position]s of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [the company should be uniquely identifiable] Who is the [Person Position] of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement? [if one, and the company should be uniquely identifiable] 1 3 1 0
@@
Who are both the [Person Position 1] and [Person Position 2] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [the company should be uniquely identifiable] Who are the [Person Position]s of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [the company should be uniquely identifiable] Who is both the [Person Position 1] and [Person Position 2] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] 1 3 1 0
Who are both the [Person Position 1] and [Person Position 2] of the company associated with [Location]? Who are the [Person Position]s of the company associated with [Location]? Who is both the [Person Position 1] and [Person Position 2] of the company associated with [Location]? [if one] 1 3 1 0
@@
Who are both the [Person Position 1] and [Person Position 2] of the company associated where [Person Name] is employed? Who are the [Person Position]s of the company associated where [Person Name] is employed? Who is both the [Person Position 1] and [Person Position 2] of the company associated where [Person Name] is employed? [if one] 1 3 1 0
@@
Who is the [Person Position 1] but not [Person Position 2] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] Who is both the [Person Position 1] and [Person Position 2] of the company which is the [Org Role(-s) (+ Sub-Role(-s))] in the agreement? [if one, and the company should be uniquely identifiable] 0 3 2 2
Who is the [Person Position 1] but not [Person Position 2] of the company associated with [Location]? [if one] Who is both the [Person Position 1] and [Person Position 2] of the company associated with [Location]? [if one] 0 3 2 1
@@
Who is the [Person Position 1] but not [Person Position 2] of the company associated where [Person Name] is employed? [if one] Who is both the [Person Position 1] and [Person Position 2] of the company associated where [Person Name] is employed? [if one] 0 3 2 3
@@
Who is the [Person Position] of the company which is the [Org Role(-s) (+ Sub-Role(-s)) 1] but not the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement? [if one, and the company should be uniquely identifiable] Who is the [Person Position] of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement? [if one, and the company should be uniquely identifiable] 0 3 2 508
What is the [Location Type] office of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] but not the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement? [if one, and the company should be uniquely identifiable] What is the [Location Type] office of the company which is both the [Org Role(-s) (+ Sub-Role(-s)) 1] and the [Org Role(-s) (+ Sub-Role(-s)) 2] in the agreement? [if one, and the company should be uniquely identifiable] 0 3 2 13